# Calibration Attention: Instance-wise Temperature Scaling for Vision Transformers

### v1

@@ 3.1 Experimental set-up
Table 1: $\downarrow$ ECE (%) for methods both before and after applying temperature scaling (Bin = 15)  [^3] We follow the experiment setup of Mukhoti et al. [2020a], utilizing their public code Mukhoti et al. [2020b]. We performed cross-validation to determine the optimal $\gamma$ for each experiment. Optimal temperature scaling values determined by validation set (included in brackets) close to 1.0 indicate well-calibrated predictions requiring minimal adjustment. For CE + BS, we follow the combination of Cross-Entropy (CE)Guo et al. [2017] and Brier Score (BS) Brier [1950a], hence the dual citation..
Dataset Model Weight Decay Guo et al. [2017] Brier Loss Brier [1950a] MMCE Kumar et al. [2018] Label Smooth Szegedy et al. [2016] Focal Loss - 53 Mukhoti et al. [2020a] Dual Focal Tao et al. [2023] CE + BS Guo et al. [2017], Brier [1950b]
Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T Pre T Post T
@@
DeiTsmall 7.46 3.54(1.10) 2.17 2.17(1.00) 7.88 2.55(1.20) 1.75 1.75(1.00) 3.27 3.27(1.00) 1.93 1.93(1.00) 5.46 3.86(1.40)
DeiTsmall+CalAttn(Ours) 6.87 1.48(1.20) 1.98 2.19(1.10) 6.51 1.86(1.20) 1.55 1.55(1.00) 3.15 3.15(1.00) 3.30 2.19(1.10) 2.88 0.86(1.10)
@@
Swinsmall 1.25 1.91(0.90) 1.12 1.12(1.00) 1.73 2.31(0.90) 6.71 1.30(0.80) 15.70 2.32(0.60) 8.54 2.40(0.70) 2.01 1.47(1.20)
Swinsmall+CalAttn(Ours) 1.74 1.74(1.00) 1.08 1.08(1.00) 1.00 1.00(1.00) 6.09 1.45(0.80) 14.97 1.66(0.60) 7.50 2.32(0.80) 0.94 0.94(1.00)
@@
Swinsmall 3.64 3.21(1.20) 1.32 0.67(0.90) 1.26 0.67(0.90) 8.09 0.56(0.60) 7.09 1.28(0.50) 5.04 0.47(0.60) 0.46 0.46(1.00)
Swinsmall+CalAttn(Ours) 0.66 0.56(1.10) 0.85 0.45(0.90) 1.16 0.51(0.0.7) 6.89 0.46(0.6) 11.75 0.84(0.50) 8.78 0.49(0.50) 0.85 0.42(0.90)
@@
DeiTsmall 13.04 2.13(1.10) 2.70 1.79(1.10) 13.66 1.25(1.10) 1.64 1.64(1.00) 10.28 2.15(1.50) 9.37 1.79(1.30) 1.69 1.55(1.10)
DeiTsmall+CalAttn(Ours) 2.37 1.38(1.40) 1.09 0.19(1.20) 2.83 1.26(1.50) 2.36 1.20(1.10) 2.15 1.11(1.30) 1.79 0.87(1.30) 0.95 0.83(1.20)
@@
Datasets. CIFAR-10 and CIFAR-100 ($50\mathrm{k}/10\mathrm{k}$ images, $32^{2}\,\mathrm{px}$) serve as high–variance, limited-data settings where calibration is notoriously difficult. To widen the difficulty spectrum we also evaluate on 200-class subset of ImageNet(Tiny) and MNIST.
Architectures. CNNs: ResNet-$50/110$, Wide-ResNet-26-10, and DenseNet-121 ($\approx 4\mathrm{M}$$23\mathrm{M}$ parameters). Transformers: ViT224[Dosovitskiy et al., 2021], DeiT-Small[Touvron et al., 2021], and Swin-Small[Liu et al., 2021] and our method implement the same backbones with a single CalAttn head.
Training. All models are trained for 350 epochs by SGD with a weight decay of $5\times 10^{-4}$ and momentum of 0.9. Learning rate set to 0.1 for the first 150 epochs, 0.01 for the following 100 epochs, and 0.001 for the remaining epochs. We retain the original data-augmentation and optimisation hyper-parameters of each backbone. CalAttn uses $\lambda{=}0.1$ in Eq. (5) for all experiments. No extra tricks or sweeps are introduced.
Baselines. We re-implement seven popular calibration losses: Weight Decay (WD) [Guo et al., 2017], Brier Score (BS) [Brier, 1950a], MMCE [Kumar et al., 2018], Label Smoothing (LS) [Szegedy et al., 2016], FocalLoss-53 [Mukhoti et al., 2020a], Dual Focal Loss (DFL) [Tao et al., 2023], and CE + BS [Guo et al., 2017, Brier, 1950b]. Every model is finally tuned with the optimal post-hoc temperature $T^{\star}$ on a $5\,\%$ validation split, following the concept of temperature scaling from Guo et al. [2017], but using the grid search procedure from Mukhoti et al. [2020a]. Specifically, $T$ is selected via a grid search over $T\in[0,0.1,0.2,\ldots,10]$ on the validation set, based on the best post-temperature-scaling ECE. We follow the random seed setting from prior public SOTA works [Wang et al., 2021, Kumar et al., 2018, Mukhoti et al., 2020a, Szegedy et al., 2016, Tao et al., 2023].
Table 2: $\downarrow$ smCE (%) for methods both before and after applying temperature scaling (Bin = 15)  [^4] The smCE metrics is based on the approach presented by Błasiok and Nakkiran [2023], with implementation publicly on GitHub Błasiok, Jarosław and Nakkiran, Preetum [2024]..

### v2

@@ 3 Experiments
This section provides a comprehensive evaluation of CalAttn across architectures, datasets, and calibration criteria  [^1] Code for the experiments is available at https://anonymous.4open.science/r/CalibrationAttention-CalAttn--D2B8.. We first report in-distribution calibration and accuracy on CIFAR-10/100, MNIST, and Tiny-ImageNet using ViT, DeiT, and Swin backbones, including results both before and after post-hoc temperature scaling. We then analyze reliability diagrams and high-confidence errors to assess safety-critical behavior. Next, we study robustness under distribution shift. We further evaluate scalability on ImageNet-1K and conduct extensive ablations on calibration head design, $\lambda$ trade-off parameter, and choice of representation used to predict temperatures. Finally, we compare CalAttn with post-hoc instance-wise methods to quantify the benefits of learning calibration.
@@ 3.1 Experimental Set-up
Datasets.
We consider CIFAR-10/100 (50k/10k, $32\times 32$), MNIST, Tiny-ImageNet, and ImageNet-1K, spanning increasing class counts and ambiguity regimes.
Architectures.
We evaluate transformer backbones ViT224 Dosovitskiy et al. (2021), DeiT-S Touvron et al. (2021), and Swin-S Liu et al. (2021). Unless explicitly stated, CalAttn augments each transformer with a single temperature head while keeping the backbone and classifier head unchanged. (We additionally include CNN baselines where relevant for comparison to prior instance-wise calibration baselines.)
Training protocol.
Unless stated otherwise, models are trained for 350 epochs with SGD (momentum 0.9, weight decay $5\times 10^{-4}$) and a step learning-rate schedule: 0.1 for first 150 epochs, 0.01 for next 100 epochs, and 0.001 for final 100 epochs. We use standard augmentation and optimization settings recommended for each backbone. CalAttn uses $\lambda=0.1$ in Eq. (3) throughout and we do not tune $\lambda$ per dataset.
Baselines.
[…]
Post-hoc temperature scaling (TS).
For each trained model, TS fits a single scalar temperature $T^{\star}$ on a held-out validation split (5% of the training data). Following Mukhoti et al. (2020a), we select $T^{\star}\in\{0.1,0.2,\ldots,10.0\}$ by grid search to minimize validation ECE (15 equal-width bins).
Table 1: $\downarrow$ ECE (%) before and after applying global temperature scaling (15 bins)  [^2] Temperature scaling fits a single scalar $T^{\star}$ on a held-out validation split (5% of training data) by grid search over $T\in\{0.1,0.2,\ldots,10.0\}$ to minimize validation ECE (15 equal-width bins), following Mukhoti et al. (2020a) and utilizing their public code Mukhoti et al. (2020b). Parentheses in Post-T report $T^{\star}$..
Dataset Model Weight Decay Guo et al. (2017) Brier Loss Brier (1950) MMCE Kumar et al. (2018) Label Smooth Szegedy et al. (2016) Focal Loss - 53 Mukhoti et al. (2020a) Dual Focal Tao et al. (2023) CE + $\lambda$ BS ($\lambda=0.1$)
Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T) Pre T Post-T (T)
@@
DeiTsmall 7.46 3.54(1.10) 2.17 2.17(1.00) 7.88 2.55(1.20) 1.75 1.75(1.00) 3.27 3.27(1.00) 3.30 2.19(1.10) 5.46 3.86(1.40)
DeiTsmall+CalAttn(Ours) 6.87 1.48(1.20) 1.98 2.19(1.10) 6.51 1.86(1.20) 1.55 1.55(1.00) 3.15 3.15(1.00) 1.93 1.93(1.00) 2.88 0.86(1.10)
@@
Swinsmall 1.95 1.91(0.90) 1.12 1.12(1.00) 1.73 2.31(0.90) 6.71 1.45(0.80) 15.70 2.32(0.60) 8.54 2.40(0.70) 2.01 1.47(1.20)
Swinsmall+CalAttn(Ours) 1.74 1.74(1.00) 1.08 1.08(1.00) 1.00 1.00(1.00) 6.09 1.30(0.80) 14.97 1.66(0.60) 7.50 2.32(0.80) 0.94 0.94(1.00)
@@
Swinsmall 3.64 3.21(1.20) 1.32 0.67(0.90) 1.26 0.67(0.90) 8.09 0.56(0.60) 7.09 1.28(0.50) 8.78 0.49(0.50) 1.86 1.86(1.00)
Swinsmall+CalAttn(Ours) 0.66 0.56(1.10) 0.85 0.45(0.90) 1.16 0.51(0.70) 6.89 0.46(0.60) 11.75 0.84(0.50) 5.04 0.47(0.60) 0.85 0.42(0.90)
@@
DeiTsmall 13.04 2.13(1.10) 2.70 1.79(1.10) 13.66 1.26(1.10) 2.36 1.64(1.30) 10.28 2.15(1.50) 9.37 1.79(1.30) 1.69 1.55(1.10)
DeiTsmall+CalAttn(Ours) 2.37 1.38(1.40) 1.09 0.19(1.20) 2.83 1.24(1.50) 1.64 1.20(1.10) 2.15 1.11(1.30) 1.79 0.87(1.30) 0.95 0.83(1.20)
@@
3.2 Main Results on ECE: before and after TS
Table 2 reports ECE before and after applying temperature scaling (TS) and CalAttn yields consistent reductions(mostly negative $\Delta$% in Fig. 3). We highlight three findings.
(1) Calibration should be considered jointly together with accuracy.
A perfectly calibrated yet inaccurate model is undesirable. We therefore report Top-1 accuracy in Table 2, and NLL alongside calibration metrics in Table 16. CalAttn improves calibration without degrading accuracy; the accuracy change is typically within a small range.
[…]
Backbone CIFAR-10 CIFAR-100
Baseline + CalAttn Baseline + CalAttn
ViT224 66.1 72.8 59.2 63.6
DeiTsmall 69.4 76.1 63.9 67.5
Swinsmall 74.5 75.8 66.7 68.2
(2) Representation-conditioned scaling yields consistent gains on transformers.
With $0.1\%$ additional parameters (Table 13), CalAttn reduces ECE for ViT/DeiT/Swin across datasets. The gain often remains after TS, indicating that CalAttn improves the underlying confidence structure rather than merely shifting logits by a global scalar.
(3) High-confidence error analysis.
Beyond ECE, we evaluate high-confidence false positives at threshold 0.90 (HCFP@0.90) and AUROC (Table 3). CalAttn consistently reduces high-confidence mistakes, demonstrating that it mitigates the most safety-critical failure mode (over-confident errors) rather than only improving calibration metrics.
Table 3: Effect of CalAttn on high-confidence errors on CIFAR-100 after post-hoc temperature scaling. $\Delta$ (%) is relative change from CE+BS; negative is improvement for error/confidence, positive is improvement for AUROC.
Backbone Metric CE+BS +CalAttn $\Delta$ (%)
ViT-224 HCFP@0.90 $\downarrow$ 21.99 17.99 18.2-18.2
Mean Conf. $\downarrow$ 97.47 92.60 5.0-5.0
AUROC $\uparrow$ 72.26 72.47 +0.3+0.3
DeiT-S HCFP@0.90 $\downarrow$ 11.99 6.99 41.7-41.7
Mean Conf. $\downarrow$ 96.27 93.50 2.9-2.9
AUROC $\uparrow$ 77.55 78.40 +1.1+1.1
Swin-S HCFP@0.90 $\downarrow$ 23.99 6.99 70.9-70.9
Mean Conf. $\downarrow$ 94.15 92.64 1.6-1.6
AUROC $\uparrow$ 73.71 81.94 +11.2+11.2
Sensitivity to $\lambda$.
[…]
3.3 Fine-grained Calibration
ECE with fixed-width binning can obscure localized miscalibration when confidence is skewed or when bins are unevenly populated. We therefore also report smECE (Table 12, Fig. 6), AdaECE (Table 17, Fig. 7), and classwise-ECE (Table 18) in Appendix F, which address complementary issues such as unequal bin occupancy and class imbalance. CalAttn maintains consistent improvements, providing evidence that gains are not an artifact of a specific binning choice.
Figure 3: ECE(%) on CIFAR-10/100 across ViT-224, DeiT-S, and Swin-S with relative changes ($\Delta$ %, “++” increase, “-” decrease).
3.4 Temporal behavior of CalAttn
Figure 4 tracks the predicted scale $s(x)=s(z)$ during training and reveals three qualitative regimes. CalAttn first recovers a near-global rescaling and then learns representation-conditioned deviations correlate with improved calibration.
Phase I — global alignment (epochs 0–30). The mean $s(z)$ rapidly decreases while the coefficient of variation $\mathrm{CV}_{s}$ drops, suggesting that CalAttn first behaves similarly to a near-global temperature rescaling.
Phase II — near-uniform modulation (epochs 30–150). After the first learning-rate decay, $\mathrm{CV}_{s}$ remains low and most samples receive similar scales, indicating limited instance-wise differentiation at this stage.
Phase III — heteroscedastic modulation (epochs 150–350). As training progresses, $\mathrm{CV}_{s}$ increases and CalAttn assigns more diverse scales across samples. Meanwhile, the mean scale crosses $s=1$, transitioning from cooling ($s>1$) to sharpening ($s<1$). This coincides with improved calibration, consistent with the hypothesis that representation-conditioned scaling becomes more expressive late in training.
Figure 4: Temporal dynamics of CalAttn on ViT-224 (CIFAR-10). Top: mean effective temperature $\tau(x)=1/s(x)$ (red) and its coefficient of variation $\mathrm{CV}_{\tau}$ (blue) over training. The dashed line marks the optimal global temperature $T^{\star}$ obtained by post-hoc temperature scaling. Bottom: kernel density estimates of $\tau(x)$ and $\lVert z_{\mathrm{CLS}}\rVert_{2}$ at three epochs.
Table 4: OOD robustness measured by AUROC (%) $\uparrow$ for detecting distribution shift from CIFAR-10 (ID) to SVHN and CIFAR-10-C.